# Floquet breathers in a modulated nonlinear lattice

### v2

@@ VII Dependence of the breather stability with the frequency
In this section, we observe the changes of stability and instability when both the breather frequency $\omega_{\text{b}}$ and $\Omega$ change simultaneously as they are commensurate as deduced before in Sec. IV. The number of possible cases, i.e, depending of the fraction $\omega_{\text{b}}/\Omega$ is huge, and the stability changes for each case are difficult to analyze. A panorama of the many cases is shown in Fig. 12, where a large number of cases for $\omega_{\text{b}}=(m_{n}/m_{m})\Omega$, with $m_{b}=10$ and $m_{m}=[2,\dots,30]$ integers are presented indicating their stability or instability type. Note that those integers are not necessarily the same as in Lemma 1 and 2 in Sec. IV.A, but their ratio it is the same. The exact value of $m_{b}$ and $m_{m}$ requires finding the fundamental frequency of the breather, which it is not difficult, but it is not the object of this work.
@@
Each breather calculation is represented by a circle with a color coding its instability as indicated by the legend. The breather circles make apparent lines with constant slope $\omega_{\text{b}}/\Omega=m_{b}/m_{n}$. A particularly interesting case is Floquet breathers with $\omega_{\text{b}}=\Omega$, where it can be seen that they can be continued into the phonon band. Other case of interest are period-doubled breathers, $\omega_{\text{b}}=\Omega/2$, for which there is a resonance with the negative phonon band $\omega_{q}^{(-)}=-\omega_{q}=-\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q))}$, which is equivalent in a system with modulation frequency $\Omega$ to $\Omega-\omega_{q}$ and the two phonon band collide. As there is no dissipation the phonons take energy from the driving, grow and become nonlinear with their frequencies separating from the unmodulated phonon band.
@@
When observing that a breather is unstable that may be due to interaction with phonons and other causes. A frequent one is that there is another breather with different structure which is stable for the same parameters. We tested for the existence and stability of breathers in quadrature, that is with a phase difference of $90^{\circ}$ with the modulating term, with interesting results. Often, but by no means always, there is an interchange of stability between the in-phase breathers and the in-quadrature breathers as can be seen in Fig. 12.
@@
In this section we concentrate into three cases of interest for ST breathers (i) Floquet breathers, with $\omega_{\text{b}}=\Omega$, period-doubled breathers $\omega_{\text{b}}=\Omega/2$, and period-halving breathers $\omega_{\text{b}}=2\Omega$. All of them have been found experimentally in photonic systems Mukherjee and Rechtsman (2023, 2020); Kang et al. (2025) and the comparison may be of interest.
@@
In order present the stability evolution, a very useful method is the simultaneous representation of the moduli and arguments of the Floquet multipliers $|\lambda_{i}|$ and $\arg{\lambda_{i}}$. Usually, the breather appear as a pair of isolated eigenvalues and the phonons and a group or band of eigenvalues, and of course, this simple picture can change a lot. We think it is more practical to defer most of the explanation to the figure captions and give here some brief indications.
@@
For Floquet breathers, the in-phase breather is unstable and the in-quadrature breather is stable, although the second one becomes unstable when its frequency collides with the phonon band as shown in Fig. 13.
@@ VIII Conclusions
We have modified a previous system corresponding to a physical model of cantilever arrays with on-site potential provide by electromagnets so as the on-site potential can be modulated both in time and space with DC and AC currents.
We have deduced their equations both in the linear and nonlinear range. We have obtained theoretically and observed numerically the change in the dispersion bands for space and time modulation separately and together. We have adapted the theory of exact breathers to time-modulated systems deducing that the modulation and breather frequencies in the moving frame should be commensurate as a necessary condition for breather existence.
We have analyzed the peculiarities of the stability analysis for time-modulated systems with the construction of and extended symplectic system.
First, breathers with a frequency multiple of the modulating one are studied and their properties analyzed.
For the modulating frequency smaller than the breather’s one, we have found single breathers and double breathers, both symmetric and anti-symmetric. When the modulating frequency becomes small a phonobreather with extended background appears.
We have analyzed with detail the evolution of the stability or instability as the frequency changes for in-phase breathers and in-quadrature breathers, observing that for some cases they interchange their stability. We have paid particular attention to equal-period, doubled-period and halved-period Floquet breathers, that appear in photonic systems, observing the evolution of the Floquet eigenvalues with the frequency, obtaining insight of the phenomena.
Some theoretical results on moving breathers in modulated systems presented in this article will be numerically studied and checked in a forthcoming article.
This study reveals that this relatively simple system provides a huge variety of Floquet breathers which may also appears in other space-time modulated materials also known as dynamic metamaterialsWu et al. (2021).
@@ Acknowledgements
All the authors would like to acknowledge Serge Aubry for the inspiration of his work. JFRA, in particular, although never worked directly with him, he did so with close collaborators as Robert Mackay and José Luis Marín in Cambridge, in 1997, where discussions with them over Aubry’s publications became his introduction to research in this field.
@@
Appendix A Notation
@@
To present formally the theory, we will use the ket and bra notation. The nice thing of that notation is that the properties of operations with vectors, change of basis, and operators appear as the simple juxtaposition of kets and bras. We present here this notation, more details can be read in Ref. Griffiths (2024) and in Ref. Archilla et al. (2003) applied to breather theory.
@@
The bra $\langle\phi|$ is the Hermitian conjugate, so $\langle\phi|$ is a row matrix with elements the complex conjugates of the elements of $|\phi\rangle$. In this way, a bracket $\langle\phi_{1}|\phi_{2}\rangle$ is the scalar product $\sum_{n=0}^{N-1}\phi_{1,n}^{*}\phi_{2,n}$. Trivially $\langle n|n^{\prime}\rangle=\delta_{n,n^{\prime}}$. So the basis $\{|n\rangle\}$ is orthonormal. $\langle n|\phi\rangle=\phi_{n}$ as the scalar product by a unit basis vector and therefore the component of $|\phi\rangle$ in the direction of the unitary vector, or the component in that basis.
@@
$=\sum_{k=0}^{N-1}\frac{1}{\sqrt{N}}\exp(-\textrm{i}\frac{\displaystyle 2\pi nk}{\displaystyle N})F_{k}(u)\,$, ($2$)
@@
A.1 Functions of position and time
@@
Appendix B Deduction of the effect of space-time modulation on the linear dispersion
In this section we consider the linear system described above in (7) which is valid for small displacements $u_{n}$.
For $\delta=0$, the system is a well known, Klein-Gordon system, having a basis of solutions $\exp(\textrm{i}qn-\omega t)$, with $\omega_{q}^{2}=\omega_{0}^{2}+4\kappa\sin^{2}(q/2)$. Note that for a given $q$ there are two frequencies $\pm\omega$ that are also frequencies for q-q.
For a finite system with $N$ units and periodic boundary conditions the possible values of $q$ are $q_{m}=2\pi m/N$, with $m=0,\dots,N-1$, this includes also negative values of the wavevector $q=-q^{\prime}$, with $q^{\prime}>0$, as any wave number $q$ is equivalent to $q\pm 2\pi n$. Also, any set of numbers $m^{\prime}=m+n$, where $n$ is any integer lead to an equivalent basis, a convenient one, centered around $q=0$ is $N/2,\dots,N/2-1$ if $N$ is even or $(N-1)/2,\dots,(N-1)/2$ if $N$ is odd. An alternative basis in real space is given[$\in[-0.5,0.5]$by $\cos(qn-\omega t)$ and $\sin(qn-\omega t)$, but the exponential basis functions a,rt=5000t=5000, for example, when the system is supposed to be near or at ore convenient analytically.
We first describe the effect of introducing the modulation term with $\delta>0$ small as observed in numerical simulations. Then we will obtain approximately the new dispersion relations for several cases.
B.1 Obtention of the numerical dispersion curves
We obtain numerically the dispersion bands with the following procedure: First we consider a relatively large system, with, for example, $N=128$ oscillators, and use as initial conditions a random distribution of positions for the variables $u_{n}$ in [-0.5,0.5], while the velocities are kept at zero. Equally, the random perturbation can be added to the momenta or both the momenta and coordinates without changing the outcome. Second, we leave the linear system (7) to evolve a relatively long time, $t=5000$, for example, when the system is supposed to be near or at thermal equilibrium and therefore all the linear modes or phonons are in principle excited. Then, we use the final positions and velocities, to run a simulation of similar length. Then, we use the results to perform the two dimensional fast fourier transform XTFFT on the variable $u_{n}(t)$, that is,
@@ B.1 Obtention of the numerical dispersion curves
with $t_{l}=\frac{\displaystyle l}{\displaystyle N_{t}}t_{f}$, $\omega_{r}=\frac{\displaystyle r}{\displaystyle N_{t}}\frac{\displaystyle 2\pi}{\displaystyle\Delta t}$, for $l,r=0\dots N_{t}-1$ and $q_{k}=\frac{\displaystyle k}{\displaystyle N}2\pi$, for $k=0\dots N-1$. The sampling interval $\Delta t=\frac{\displaystyle t_{f}}{\displaystyle N_{t}}$ is larger that the integration step to avoid very large matrices as the maximum frequency $2\pi/\Delta t$ does not need to be very large because the frequencies are of the order or $\omega_{0}$ and $\Omega$. On the other hand the frequency precision $\Delta\omega=\frac{\displaystyle 2\pi}{\displaystyle N_{t}\Delta t}=\frac{\displaystyle 2\pi}{\displaystyle t_{f}}$, is inversely proportional to $t_{f}$, therefore a large value of it is convenient, both to obtain precision for the frequency and for the results to be generic enough.
@@
The intensities are defined as $I(\omega_{i},q_{n})=|F(\omega_{i},q_{n})|^{2}$, normalized to unity. To visualize the dispersion curves, we represent a number of lines in contour plots, typically ten, to get a clear picture. However, all the modes do not have the same probability to occur, the density of states may be small, and theoretically existing modes have small intensities and are difficult to discriminate from the noise background, those dispersion lines are weak. Presently, we do not have a means to calculate theoretically those intensities.
@@
This method is changed to check that the results do not depends on the details, the random distribution can be applied only to momenta or to both coordinates and momenta; we can change the thermalization time, simulation time, sampling time interval, number of particles, etc. This method will also be used with a superposed breather in other sections.
@@
The result of the simulation after thermalization are also used to obtain energy density plots. An example of both is shown in Fig. 2.
We can see that new dispersion bands appear roughly parallel to the unmodulated one and with their maxima and minima displaced.
B.2 Dispersion bands modification with only space modulation

### v3

@@ VII Dependence of the breather stability with the frequency
In this section, we observe the changes of stability and instability when both the breather frequency $\omega_{\text{b}}$ and $\Omega$ change simultaneously as they are commensurate as deduced before in Sec. IV. The number of possible cases, i.e, depending of the fraction $\omega_{\text{b}}/\Omega$ is huge, and the stability changes for each case are difficult to analyze. A panorama of the many cases is shown in Fig. 12, where a large number of cases for $\omega_{\text{b}}=(m_{\text{b}}/m_{\text{m}})\Omega$, with $m_{\text{b}}=10$ and $m_{\text{m}}=[2,\dots,30]$ integers are presented indicating their stability or instability type. Note that those integers are not necessarily the same as in Lemma 1 and 2 in Sec. IV.A, but their ratio is the same.
@@
Each breather calculation is represented by a circle with a color coding its instability as indicated by the legend. The breather circles make apparent lines with constant slope $\omega_{\text{b}}/\Omega=m_{\text{b}}/m_{\text{n}}$. A particularly interesting case are equal-period breathers with $\omega_{\text{b}}=\Omega$, where it can be seen that they can be continued into the phonon band. Other case of interest are period-doubled breathers, $\omega_{\text{b}}=\Omega/2$, for which there is a resonance with the negative phonon band $\omega_{q}^{(-)}=-\omega_{q}=-\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q))}$, which is equivalent in a system with modulation frequency $\Omega$ to $\Omega-\omega_{q}$ and the two phonon bands collide. As there is no dissipation the phonons take energy from the driving, grow and become nonlinear with their frequencies separating from the unmodulated phonon band.
@@
When observing that a breather is unstable, that may be due to interaction with phonons and other causes. A frequent one is that there is another breather with different structure which is stable for the same parameters. We tested for the existence and stability of breathers in quadrature, that is, with a phase difference of $\pi/2$ with the modulating term, with interesting results. Often, but by no means always, there is an interchange of stability between the in-phase breathers and the in-quadrature breathers as can be seen in Fig. 12.
@@
In this section we concentrate into three cases of interest for ST breathers (i) equal-period breathers, with $\omega_{\text{b}}=\Omega$, period-doubled breathers $\omega_{\text{b}}=\Omega/2$, and period-halving breathers $\omega_{\text{b}}=2\Omega$. All of them have been found experimentally in photonic systems Mukherjee and Rechtsman (2023, 2020); Kang et al. (2025) and the comparison may be of interest.
@@
In order to present the stability evolution, a very useful method is the simultaneous representation of the moduli and arguments of the Floquet multipliers $|\lambda_{i}|$ and $\arg{\lambda_{i}}$. Usually, the the existence of a breather appears as a pair of isolated eigenvalues corresponding to internal modes of the breather, and the phonons as a group or band of eigenvalues, and, of course, this simple picture can change a lot. We think it is more practical to defer most of the explanation to the figure captions and give here some brief indications.
@@
For equal-period Floquet breathers, the in-phase breather is unstable and the in-quadrature breather is stable, although the second one becomes unstable when its frequency collides with the phonon band as shown in Fig. 13.
@@ VIII Conclusions
In this work we have studied the existence and stability of Floquet breathers in a space-time modulated nonlinear lattice. The physical system under study is a variation of a previously presented design of a cantilever array, with an on-site potential provided by magnetic forces created by electromagnets, which are now driven by DC and AC currents to create a modulation both in time and space.
The nonlinear dynamical equations are derived, and is linear limit discussed. The dispersion bands, and its changes due to space, time, and space-time modulation have been obtained analytically and numerically.
We have adapted the theory of exact breathers to time-modulated systems, concluding that the modulation and breather frequencies in the moving frame should be commensurate as a necessary condition for breather existence.
The peculiarities of the stability analysis for time-modulated systems have been also analyzed, with the construction of an extended symplectic system.
First, breathers with a frequency multiple of the modulating one are studied and their properties analyzed. For modulating frequencies smaller than the breather frequency, single breathers and double breathers are found, both symmetric and anti-symmetric. When the modulating frequency becomes small a phonobreather with extended background appears.
The evolution of the stability or instability as the frequency changes is analyzed in detail for in-phase and in-quadrature breathers, observing that for some cases they interchange their stability. Attention was paid to equal-period, doubled-period and halved-period Floquet breathers, that appear in photonic systems, observing the evolution of the Floquet eigenvalues with the frequency, obtaining insight of the phenomena.
Some theoretical results on moving breathers in modulated systems presented in this article will be numerically studied and checked in a forthcoming article. This study reveals that this relatively simple system provides a huge variety of Floquet breathers which may also appears in other space-time modulated materials also known as dynamic metamaterials.
@@ Acknowledgements
All the authors would like to acknowledge Serge Aubry for the inspiration of his work. JFRA, in particular, although never worked directly with him, he did so with close collaborators as Robert Mackay and José Luis Marín in Cambridge, in 1997, where discussions with them over common publications with Aubry became his introduction to research in this field.
@@
Appendix A Numerical experiments, thermalization and noisy systems
Along several parts of this work and in the sections below, we use the terms thermalization, near thermalization, or noisy systems. In this appendix we specify what we understand by those related concepts.
We perform numerical simulations and analytical calculations to obtain properties of the system as the initial coordinates and momenta for breathers, or the discrete Fourier components, or to find their stability properties. These simulations are supposed to produce a definite and precise outcome. The imprecission is due to shortcomings of the numerical method, the parameters chosen, as time integration steps, the number of Fourier components and so on.
We also perform numerical experiments to observe the system properties, similarly to physical experiments. The initial conditions are chosen at random with some desired energy or averages, and we observe the system evolve and its route to an approximate thermal equilibrium, and its behavior once this thermal equilibrium is attained. An important example is to obtain the numerically experimental dispersion relation with different modulations as explained in Appendix C.
The term approximate appears because a rigorous definition of thermal equilibrium consists of the property that the statistical ensemble compatible with the macroscopic constrains allows the system to realize all the microscopic states with equal probability.Reif (2009) This is not achievable in practice, and the use of the ergodic theorem that a single realization of the ensemble also occupy all the microstates with equal probability if allowed infinite time is also difficult in practice.
Therefore, we consider that the system is in thermal equilibrium when some of its consequences are approximately met. We understand that (i) the system forgets its microscopic initial conditions, i.e., positions and velocities but not energy and other quantities; (ii) its local energy probability distribution approaches the Boltzmann distribution; (iii) identical particles have the same properties over time; (iv) the participation function aproaches $N/2$ (see below); (v) the site of the particle with peak local energy moves along the lattice, perhaps with space-time correlation of some time when fluctuations bring about wave groups with some lifetime (See. Fig. 2) but not permanently unless they are a consequence of the system properties. Some temporary correlations are also properties of the system as phonon wave groups. Those are detected to construct, for example, the numerically experimental dispersion relations as explained in App. C.1.
This subject is dealt in Ref. Archilla et al., 2025 in the following way. Let us consider a system of particles with weak coupling in the microcanonical ensemble, we can suppose that each particle is in contact with a thermal bath corresponding to the rest of the system at a given temperature $T$, or thermodynamic $\beta=1/\mu$, with $\mu=kT$, $k$ being the Boltzmann constant. Within this approximation, the probability of a particle having an energy $e_{n}$ is proportional to $\exp(-\beta e_{n})$,Reif (2009) with $\mu$ equal to twice the average kinetic energy, according to the virial theorem.
Figure 15: (Top:) Path to thermal equilibrium for a space-modulated system with $h=2\pi/3$. The initial conditions are a random distribution of velocities with the coordinates at rest. The theoretical value at equilibrium is $P=N/2$ (in red), but with large fluctuations for a $N=120$ system. The logarithmic scale for time allows the observation both at the short scale at the beginning but also for longer times; (Bottom:) Numerically experimental distribution function (blue) compared with the theoretical one (red). The first bin is discarded in the experimental curve. Parameters: $N=120$, $\kappa=0.10$, $\delta=0.05$, $\Omega=0$, $h=2\pi/3$.
Imposing the probability one for a particles to have all the energies from 0 to $\infty$, the probability density of a particle having energy $e_{n}$ is $\rho(e_{n})=\exp(-e_{n}/\mu)/\mu$, also known as the exponential distribution. The average particle energy is $\langle e_{n}\rangle=\mu$, with standard deviation also $\mu$. Averages are in principle for the statistical ensemble, but due to the ergodic theorem, the average can be made in time and also at the different identical particles.
Figure 16: (Top:) Path to thermal equilibrium for the particles outside a breather core. Initial conditions are a random distribution of velocities with the coordinates at rest except for the initial breather coordinates at a core of $1$ particles in a system with $N=160$ particles. The theoretical value at equilibrium is $P=N/2$ (in red), but with large fluctuations for a N=16011=149N=160-11=149 system. In this case, a shorter time than in Fig. 15 is shown in linear scale. (Bottom:) Numerical experimental distribution function (blue) compared with the theoretical one (red) outside the breather core. Parameters: $N=160$, $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\Omega=\omega_{\text{b}}/20=0.049$, $h=0$.
An easily measurable consequence of this distribution is the behavior of the participation function $P=(\sum e_{n})^{2}/\sum e_{n}^{2}$. This is a function with value 1 if all the energy is located at a single particle and $N$ if the energy is evenly distributed among the system number of particles $N$. The value of $P$ at thermal equilibrium using the exponential distribution is $N/2$, with relatively large oscillations of $P$ due to the large fluctuations of $e_{n}$ for the exponential distribution. The amplitude of the oscillation of $P$ around $N/2$ depends on the system size, the larger the size, the relative value of the oscillation becomes smaller as the relative standard deviation is or the order of $1/\sqrt{N}$. As the distribution $\rho(e_{n})$ is not exact, the actual value of the average $P$ can be a few particles above or below $N/2$, also changing with the time observation window, due to the large fluctuations of a small system with $N\simeq 100$.
We can check these properties for our system with different types of modulation. The exponential distribution fails at low energies when the model of a particle with energy $e_{n}$ interacting with the nearest neighbors is not correct as part of the energy is precisely the interaction energy with the nearest neighbors. Often the first bin to obtain the numerical distribution have to be discarded.
As an example, let us consider the simulations to obtain the experimental dispersion relation for the system modulated in space with $h=2\pi/3$ with the theoretical and numerical results presented in Appendix C.2.
Noisy systems
The concept of noisy system appears when we study breathers. The linear stability of breathers, and its nonlinear stability for small perturbations, is analyzed in Sec. IV.D, but another complementary technique of particular physical interest is to observe the breather behavior in presence of temperature. A system with a breather is not thermalized, because eventually the breather will disappear but the breather life can be very long for some systems and breather energies.Archilla et al. (2025) A noisy system is a system that it is in an approximate thermal equilibrium, in the sense explained above, outside the breather core. The breather core is a few particles around the breather center, depending on the breather localization.
We add random velocities with some convenient kinetic energy to the initial coordinates for a system with a breather and perform numerical experiments with similar outcomes. The participation function of the whole system will be a few particles, but outside the breather core the same properties at thermal equilibrium considered above can be observed. This is illustrated in Fig. 16.
Appendix B Notation
@@
To present formally the theory, we will use the ket and bra notation. The nice thing of that notation is that the properties of operations with vectors, change of basis, and operators appear as the simple juxtaposition of kets and bras. We present here this notation, more details can be read in Ref. Griffiths, 2024 and in Ref. Archilla et al., 2003 applied to breather theory.
@@
The bra $\langle\phi|$ is the Hermitian conjugate, so $\langle\phi|$ is a row matrix with elements the complex conjugates of the elements of $|\phi\rangle$. In this way, a bracket $\langle\phi_{1}|\phi_{2}\rangle$ is the scalar product $\sum_{n=0}^{N-1}\phi_{1,n}^{*}\phi_{2,n}$. Trivially $\langle n|n^{\prime}\rangle=\delta_{n,n^{\prime}}$. So the basis $\{|n\rangle\}$ is orthonormal. $\langle n|\phi\rangle=\phi_{n}$ is the scalar product by a unit basis vector and therefore the component of $|\phi\rangle$ on that basis vector.
@@
$=\sum_{k=0}^{N-1}\frac{1}{\sqrt{N}}\exp(-\textrm{i}\frac{\displaystyle 2\pi nk}{\displaystyle N})F_{k}(u)\,$, (29)
@@
B.1 Functions of position and time
@@
Appendix C Derivation of the effect of space-time modulation on the linear dispersion
The dispersion relations obtained in numerical experiments and the theoretical ones have been already presented in Sec. III, with the exception of only space modulation presented below as it was not essential for the rest of the article. Here, after some short introduction, we describe the method to visualize the dispersion relations in numerical experiments followed by the theoretical derivations in several cases.
We consider the linear system described above in (7) which is valid for small displacements $u_{n}$.
For $\delta=0$, the system is a well known, Klein-Gordon system, having a basis of solutions $\exp(\textrm{i}qn-\omega t)$, with $\omega_{q}^{2}=\omega_{0}^{2}+2\kappa(1-\cos(q))$. Note that for a given $q$ there are two frequencies $\pm\omega$ that are also frequencies for q-q.
For a finite system with $N$ units and periodic boundary conditions the possible values of $q$ are $q_{m}=2\pi m/N$, with $m=0,\dots,N-1$, this includes also negative values of the wavevector $q=-q^{\prime}$, with $q^{\prime}>0$, as any wave number $q$ is equivalent to $q\pm 2\pi n$. Also, any set of numbers $m^{\prime}=m+n$, where $n$ is any integer lead to an equivalent basis. A convenient one, centered around $q=0$ is $N/2,\dots,N/2-1$ if $N$ is even, or $(N-1)/2,\dots,(N-1)/2$ if $N$ is odd. An alternative basis in real space is given by $\cos(qn-\omega t)$ and $\sin(qn-\omega t)$, but the exponential functions basis is more convenient analytically.
C.1 Obtention of the dispersion curves in numerical experiments
We obtain numerically what we could call the experimental dispersion curves in an approximate thermal equilibrium, as explained with detail in App. A, with the following procedure: First we consider a relatively large system, with, for example, $N=128$ oscillators, and use as initial conditions a random distribution of positions for the variables $u_{n}$ in [-0.5,0.5], Equally, the random perturbation can be added to the momenta or both the momenta and coordinates without changing the outcome. Second, we leave the linear system (7) to evolve a relatively long time, $t=5000$, for example, when the system is supposed to be near or at thermal equilibrium and therefore all the linear modes or phonons are in principle excited. We can use also the full nonlinear system but with conveniently small values of the coordinates or momenta so that the system is near the linear limit. Then, we use the final positions and velocities, to run a simulation of similar length. We use the results to perform the two dimensional discrete Fourier transform in space and time (XTDFT) on the variable $u_{n}(t)$, that is,
@@ C.1 Obtention of the dispersion curves in numerical experiments
with $t_{l}=\frac{\displaystyle l}{\displaystyle N_{t}}t_{f}$, $\omega_{r}=\frac{\displaystyle r}{\displaystyle N_{t}}\frac{\displaystyle 2\pi}{\displaystyle\Delta t}$, for $l,r=0\dots N_{t}-1$ and $q_{k}=\frac{\displaystyle k}{\displaystyle N}2\pi$, for $k=0\dots N-1$. The sampling interval $\Delta t=\frac{\displaystyle t_{f}}{\displaystyle N_{t}}$ is larger that the integration step to avoid very large matrices as the maximum frequency $2\pi/\Delta t$ does not need to be very large because the frequencies are of the order of $\omega_{0}$ and $\Omega$. On the other hand the frequency precision $\Delta\omega=\frac{\displaystyle 2\pi}{\displaystyle N_{t}\Delta t}=\frac{\displaystyle 2\pi}{\displaystyle t_{f}}$, is inversely proportional to $t_{f}$, therefore a large value of it is convenient, both to obtain precision for the frequency and for the results to be generic enough.
@@
The intensities are defined as $I(\omega_{i},q_{n})=|F(\omega_{i},q_{n})|^{2}$, normalized to unity. To visualize the dispersion curves, we represent a number of lines in contour plots, typically ten, to get a clear picture. However, all the modes do not have the same probability to occur, the density of states may be small, and theoretically existing modes may have small intensities and are difficult to discriminate from the noise background, those dispersion lines are weak. Presently, we do not have a means to calculate theoretically those intensities.
@@
This method is changed to check that the results do not depends on the details, the random distribution can be applied only to momenta or to both coordinates and momenta; we can change the thermalization time, simulation time, sampling time interval, number of particles, etc. This method can also be used with a superposed breather in other sections.
@@
The result of the simulation after thermalization are also used to obtain energy density plots. An example of both is shown in Fig. 2. We can see that new dispersion bands appear roughly parallel to the unmodulated one and with their maxima and minima displaced.
C.2 Dispersion bands modification with only space modulation